# EdgeProfiler: A Fast Profiling Framework for Lightweight LLMs on Edge Using Analytical Model

### v2

@@ VI Conclusion
In this paper we have propose EdgeProfiler, a fast profiling framework for lightweight LLMs on edge devices using an analytical model. EdgeProfiler framework reveals how hardware characteristics, model size, and numerical precision jointly shape inference performance and efficiency. We present our findings from three widely used edge devices. Across all platforms, reducing precision from FP32, FP16, and INT8 yields substantial end-to-end latency and energy savings. INT8 cuts the latency by $\sim$75and energy by $\sim$75relative to FP32. On low-power devices, Raspberry Pi, memory, and I/O bound stage often dominate, whereas on more capable hardware like Jetson Nano, compute latency becomes negligible, and I/O or network transfer surface as the new bottleneck. Building upon this foundation, future work will focus on refining profiling granularity, integrating real-world deployment feedback, and expanding support for emerging hardware platforms to accelerate practical, scalable, and energy-efficient LLM applications on the edge.

### v3

@@ VI Conclusion
In this paper, we have propose EdgeProfiler, a fast profiling framework for lightweight LLMs on edge devices using an analytical model. EdgeProfiler framework reveals how hardware characteristics, model size, and numerical precision jointly shape inference performance and efficiency. We present our findings from three widely used edge devices. Across all platforms, reducing precision from FP32, FP16, and INT8 yields substantial end-to-end latency and energy savings. On low-power devices, Raspberry Pi, memory, and I/O bound stage often dominate, whereas on more capable hardware like Jetson Nano, compute latency becomes negligible, and I/O or network transfer surface as the new bottleneck. Building upon this foundation, future work will focus on refining profiling granularity, integrating real-world deployment feedback, and expanding support for emerging hardware platforms to accelerate practical, scalable, and energy-efficient LLM applications on the edge.